# REDS: Resource-Efficient Deep Subnetworks for Dynamic Resource Constraints

### v2

@@ Appendix A Iterative Knapsack Theory
A.1 Knapsack for depth-wise convolutions
In this section we define a generalized knapsack framework for depth-wise convolutions. Given a $m\times n$ input matrix $A$, the first layer $L_{0}$ is a standard convolutional layer (cf.f Fig. 2) applied to $A$ that consists of $N_{0}$ filters and hence produces $N_{0}$ output channels. It is followed by depth-wise convolutional blocks $B_{1},\ldots,B_{d}$, where each block $i$ is built by a depth-wise layer $B_{i}^{D}$ and a point-wise layer $B_{i}^{P}$ for $i=1,\ldots,d$. A depth-wise layer applies one filter on each of the $N_{i-1}$ input channels. The point-wise layer $B_{i}^{P}$ then applies $N_{i}$ filters, where the kernel number for each filter has to be equal to the number of filters of the previous layer. If we want to reduce the size of the network, we can decide on how many filters we use at $L_{0}$ and at each $B_{i}^{P}$. E.g., if we choose $k$ filters at $L_{0}$ the layer $B_{1}^{D}$ will have $k$ filters and one filter of $B_{1}^{P}$ will have $k$ as kernels number. The resulting optimization method can be used for structural pruning of a neural network, i.e., for choosing an optimal number ($\leq N_{0}$) of filters of $L_{0}$ and (at most $N_{i}$) filters of $B_{i}^{P}$ such that we maximize performance of the network obeying a constraint on time or space complexity, which in REDS is the number of MACs.
Formally, we introduce an integer decision variable $x_{0}$ to decide on the number of filters we use at $L_{0}$, and $x_{i}$ on the number of filters to use at each $B_{i}^{P}$. Since every computational unit consumes computational resources and contributes to the overall accuracy, we also introduce decision variables that control whether a unit is used by a subnetwork or not. For $L_{0}$ we introduce $N_{0}$ binary variables $y_{1},\ldots,y_{N_{0}}$, and for every block $B_{i}$, $i=1,\ldots,d$, we introduce binary variables $f_{k}^{i}$ and $g_{kt}^{i}$, where $k\in\{1,\ldots,N_{i}\}$ and $t\in\{1,\ldots,N_{i-1}\}$, $f_{k}^{i}$ to indicate whether a filter $k$ is used in $B_{i}^{P}$, and $g_{kt}^{i}$ if filter $k$ with kernel $t$ of $B_{i}^{P}$ is used. For $B_{i}^{D}$ we introduce the decision variables $d_{t}^{i}$ to decide if a depth-wise filter $t\in\{1,\ldots,N_{i-1}\}$ is used.
In the model $P^{1}_{i}$ is the importance score of a standard convolution filter $i$ in the first layer and $W_{1}$ is its number of MACs. For each depth-wise point-wise block $i$, $P^{i}_{t}$ is the importance score for the depth-wise filter $t$, $P^{i}_{kt}$ is the importance score of the corresponding kernel $k$ of the point-wise filter $t$ in the subsequent layer $N_{i}$. Due to permutation invariance (Sec. 3.1) the importance scores $P^{i}$ and $P^{i}_{t}$ are stored in descendant order. The problem is formulated as:
$\max$ $\sum_{i=1}^{N_{0}}y_{i}\cdot P^{1}_{i}+\sum_{i=1}^{d}\sum_{t=1}^{N_{i-1}}\left(d_{t}^{i}\cdot P^{i}_{t}+\sum_{k=1}^{N_{i}}g_{kt}^{i}\cdot P^{i}_{kt}\right)\;(1)$
$s.t.$ $\sum_{i=1}^{N_{0}}y_{i}\cdot W_{1}+\sum_{i=1}^{d}\sum_{t=1}^{N_{i-1}}\left(d_{t}^{i}\cdot W_{2}+\sum_{k=1}^{N_{i}}g_{kt}^{i}\cdot W_{3}\right)\leq C\;(2)$
$\sum_{i=1}^{N_{0}}y_{i}=x_{0}\;(3)\;\;\;\;\text{and}\;\;\;\;\sum_{t=1}^{N_{i-1}}d_{t}^{i}=x_{i-1}\enspace\forall i\;(4)$
$\sum_{k=1}^{N_{i}}f_{k}^{i}=x_{i}\enspace\forall i\;(5)\;\;\;\;\text{and}\;\;\;\;f_{k}^{i}\geq f_{k+1}^{i}\enspace\forall i\;(6)$
$g_{kt}^{i}\leq f_{k}^{i}\enspace\forall i,k,t\;(7)\;\;\;\;\text{and}\;\;\;\;f_{k}^{i}\leq\sum_{t=1}^{N_{i-1}}g_{kt}^{i}\enspace\forall i,k\;(8)$
$\sum_{t=1}^{N_{i-1}}g_{kt}^{i}\leq x_{i-1}\enspace\forall i,k\;(9)$
$\sum_{t=1}^{N_{i-1}}g_{kt}^{i}\geq x_{i-1}-(1-f_{k}^{i})\cdot N_{i-1}\enspace\forall i,k\;(10)$
The knapsack for depth-wise convolutions is described as:
  • (1)
    The objective function maximizes the total importance score of the chosen architecture.
  • (2)
    Solution MACs must comply with the constraint C.
  • (3)
    The number of filters chosen in the first convolution layer.
  • (4)
    The number of filters chosen in the depth-wise layer $i$ must match the number of filters picked in the previous layer $i-1$.
  • (5)
    The number of the point-wise filters chosen in layer $i$.
  • (6)
    Point-wise filters are chosen in ascending order to impose a contiguous solution.
  • (7)
    If kernel $t$ of filter $k$ is chosen then the whole filter $k$ is chosen.
  • (8)
    A point-wise filter $k$ is chosen only if one of its kernels is chosen.
  • (9)
    The number of kernels in the filter $k$ of point-wise layer $i$ must $\leq$ the number of filters taken in the previous depth-wise layer.
  • (10)
    If filter $k$ in layer $i$ is chosen then the constraints (9) and (10) together ensure that the number of kernels $t$ of filter $k$ in layer $i$ equals the number of point-wise filters in the previous block (i.e., the number of filters in the depth-wise layer $i$). If filter $k$ in layer $i$ is not chosen, constraints (9) and (10) together imply that all kernels $t$ of filter $k$ at layer $i$ are zero (the right-hand side of (9) is $\leq 0$, since $N_{i-1}$ is an upper bound on $x_{i-1}$).
A.2 Iterative knapsack problem
@@
Here $I(Opt_{c})=\{1,2,3\}$. $D_{c/2}$ then selects one of these items and no more items fit into the knapsack. $Opt_{c/2}$ selects item $4$, which shows that the ratio of $\frac{1}{2}$ is tight. ∎
@@ Appendix B Training Details
Table 4 summarizes the hyper-parameters used to train different networks. We refer to Zhang et al. 2017 regarding the description of the network architectures adopted in this paper (referred to as DNN, CNN and DS-CNN, sizes S and L). We used TensorFlow (Abadi et al., 2016) version 2.11. All models were trained on a workstation with 16 NVIDIA Tesla K80 GPUs and 32 Intel Xeon CPUs. To conduct all our experiments and compute the baselines we trained and optimized over 100 models. This translated into $>$1000 h of compute. We always use 80:10:10 split for training, validation, and testing. All results constitute averages over three runs.
@@ C.1 REDS performance on DNN and CNN architectures
Figure 11: Analysis of the subnetwork architecture obtained by the knapsack BU heuristics. From left to right: DNN S, DNN L, CNN S and CNN L on Google Speech Commands. The patterns as to which computational units constitute a child subnetwork are architecture-specific.
@@ C.2 REDS performance with 10 nested subnetworks
Table 7 and Fig. 12 show the performance of DNN, CNN and DS-CNN on Google Speech Commands, when REDS structure comprises 10 subnetworks, compared to 4 subnetworks in the main paper. A larger number of subnetworks does not degrade model accuracy.
@@
Figure 12: REDS size S (top row) and L (bottom row) architectures analysis finetuned on Google Speech Commands Warden (2018) with ten subnetworks. The plots from left to right show the subnetworks size, the subnetworks accuracy and the subnetworks inference time as a function of MAC percentage.
C.3 REDS on FMNIST and CIFAR10
The results in Table 8 and Table 9 show REDS performance using DS-CNN architecture of size S on FMNIST and CIFAR10. BU heuristic was used to obtain the results. REDS supports a different data domain without degrading the accuracy of the pre-trained model, reported in the header row. Compared to the state-of-the-art such as $\mu$NAS Liberis et al. (2021), REDS demonstrates a faster architecture search time for both FMNIST and CIFAR10. In the former, REDS takes 19 minutes as opposed to 3 days; in the latter, REDS takes 90 minutes as opposed to 39 days while requiring less memory for model storage for both datasets. After finding and freezing the 25% MACs subnetwork architecture, the BU heuristic takes only a few seconds to find the other 50% and 75% MACs subnetworks architectures.
@@ C.3 REDS on FMNIST and CIFAR10
100% 91.6 ±$\pm 0.2$ 128.54
75% 91.51 ±$\pm 0.28$ 107.73 1.58 [s]
50% 90.75 ±$\pm 0.32$ 87.4 5.83 [s]
25% 89.22 ±$\pm 0.45$ 66.63 19 [m]
Table 8: Analysis of the BU knapsack subnetworks obtained from a depth-wise separable convolutional (DS-CNN) S network, pre-trained on FMNIST Xiao et al. (2017). REDS supports a different data domain without degrading the accuracy of the pre-trained model, reported in the header row.
@@
100% 81.07 ±$\pm 0.71$ 128.54
75% 80.17 ±$\pm 0.69$ 109.41 2.89 [s]
50% 76.72 ±$\pm 1.37$ 88.01 10.59 [s]
25% 68.66 ±$\pm 1.65$ 69.63 90 [m]
Table 9: The same evaluation as in Table 8 for CIFAR10 Krizhevsky et al. (2009).

### v3

@@ Appendix A Iterative Knapsack Theory
Here $I(Opt_{c})=\{1,2,3\}$. $D_{c/2}$ then selects one of these items and no more items fit into the knapsack. $Opt_{c/2}$ selects item 444, which shows that the ratio of $\frac{1}{2}$ is tight. ∎
@@ Appendix B Training Details
Table 7 summarizes the hyper-parameters used to train different networks. We refer to Zhang et al. (2017) regarding the description of the network architectures adopted in this paper (referred to as DNN, CNN and DS-CNN, sizes S and L).
@@ C.1 REDS performance on DNN and CNN architectures
Figure 12: Analysis of the subnetwork architecture obtained by the knapsack BU heuristics. From left to right: DNN S, DNN L, CNN S and CNN L on GSC. The patterns as to which computational units constitute a child subnetwork are architecture-specific.
@@ C.2 REDS performance with 10 nested subnetworks
Table 10 and Fig. 13 show the performance of DNN, CNN and DS-CNN on GSC, when REDS structure comprises 10 subnetworks, compared to 4 subnetworks in the main paper. A larger number of subnetworks does not degrade model accuracy.
@@
Figure 13: REDS size S (top row) and L (bottom row) architectures analysis finetuned on Google Speech Commands (Warden, 2018) with ten subnetworks. The plots from left to right show the subnetworks size, the subnetworks accuracy and the subnetworks inference time as a function of MAC percentage.
C.3 REDS on Fashion-MNIST and CIFAR10
The results in Table 11 and Table 12 show REDS performance using DS-CNN architecture of size S on Fashion-MNIST and CIFAR10. BU heuristic was used to obtain the results. REDS supports a different data domain without degrading the accuracy of the pre-trained model, reported in the header row. Compared to the state-of-the-art such as $\mu$NAS (Liberis et al., 2021), REDS demonstrates a faster architecture search time for both Fashion-MNIST and CIFAR10. In the former, REDS takes 19 minutes as opposed to 3 days; in the latter, REDS takes 90 minutes as opposed to 39 days while requiring less memory for model storage for both datasets. After finding and freezing the 25% MACs subnetwork architecture, the BU heuristic takes only a few seconds to find the other 50% and 75% MACs subnetworks architectures.
@@ C.3 REDS on Fashion-MNIST and CIFAR10
100% 91.6 ±0.2$\pm$ 0.2± 0.2 128.54
75% 91.51 ±0.28$\pm$ 0.28± 0.28 107.73 1.58 [s]
50% 90.75 ±0.32$\pm$ 0.32± 0.32 87.4 5.83 [s]
25% 89.22 ±0.45$\pm$ 0.45± 0.45 66.63 19 [m]
Table 11: Analysis of the BU knapsack subnetworks obtained from a depth-wise separable convolutional (DS-CNN) S network, pre-trained on Fashion-MNIST (Xiao et al., 2017). REDS supports a different data domain without degrading the accuracy of the pre-trained model, reported in the header row.
@@
100% 81.07 ±0.71$\pm$ 0.71± 0.71 128.54
75% 80.17 ±0.69$\pm$ 0.69± 0.69 109.41 2.89 [s]
50% 76.72 ±1.37$\pm$ 1.37± 1.37 88.01 10.59 [s]
25% 68.66 ±1.65$\pm$ 1.65± 1.65 69.63 90 [m]
Table 12: The same evaluation as in Table 11 for CIFAR10 (Krizhevsky et al., 2009).